A splinter starts from rest and uniformly accelerates at a rate of 2m/s2 for 5.0 seconds. What is his final velocity?


a. 20 m/s

b. 10 m/s

c. 5 m/s

d. 25 m/s


1
Expert's answer
2022-02-16T13:22:48-0500

Given:

"v_i=0"

"a=2\:\rm m/s^2"

"t=5.0\:\rm s"

The final velocity

"v_f=v_i+at=0+2*5.0=10\:\rm m/s"

Answer: b. 10 m/s